OpenStudy (pounce0129):

solve the system by using the substitution method. write your answer as an ordered pair. r+3s=8 s-r=-4

OpenStudy (kropot72):

r + 3s = 8 ...........(1) s - r = -4 ............(2) If you subtract 3s from both sides of equation (1) you can find r in terms of s. Then substitute this value of r into equation (2) and rearrange to find the value of s.

OpenStudy (pounce0129):

....? im not that smart

OpenStudy (kropot72):

Taking the first step we subtract 3s from both sides of equation (1) and get: r + 3s - 3s = 8 - 3s ..............(3) Now simplify the left hand side of equation (3). Can you do that?